find the equation of the ellipse whose major axis is twice its minor axis and passes through the point (0,1)
 

Grade:12

1 Answers

Deepak Kumar Shringi
5 years ago
major axis = 2 minor axis
a=2b
passes from (0,1) so
x^2/a^2+y^2/b^2=1 let the equation now put a=2b
x^2/4b^2+y^2/b^2=1
pass from (0,1)
1/b^2=1 so b=1 so a=2 equation will be
x^/4+y^2=1